# Diffusion Self-Weighted Guidance for Offline Reinforcement Learning

### v1

@@ 3.3 Augmented exact guidance
The required score to sample from the target distribution in Eq. (16) is given by
@@
The proof for Proposition 3.3 follows as an extension of cep to our setting (see Appendix A.1).
@@
Since the extended formulation in Eq. (16) is defined over $\mathbf{z}$, the gradients required to implement guidance in Eq. (18) are taken with respect to $\mathbf{z}_{k}$, rather than $\mathbf{a}_{k}$ (as in Eq. (13)). This is a desired feature, since the actions and weights are highly coupled and thus guiding them separately can lead to inefficient sampling.
The key insight of our proposal is that the intractable term in Eq. (18) can be recovered directly from the DM described above using the so-called data prediction formula efron2011tweedie ; variational_diffusion_models given by:
@@
Using Eq. (19), and since the extraction function $\phi_{w}(\cdot)$ is linear, the target score $\nabla_{\mathbf{z}_{k}}\log p_{k}(\mathbf{z}_{k})$ in Eq. (18) can be computed with the proposed DM, as stated in Theorem 3.5 (see the proof in Appendix A.2).
@@
Let $\epsilon^{*}_{\theta}(\mathbf{z}_{k},k)$ be the DM found as the global minimizer of Eq. (17). The exact target score of Eq. (18) is given by:
$\nabla_{\mathbf{z}_{k}}\log p_{k}(\mathbf{z}_{k})=\underbrace{\nabla_{\mathbf{z}_{k}}\log q_{k}(\mathbf{z}_{k})}_{\frac{-\epsilon^{*}_{\theta}(\mathbf{z}_{k},k)}{\sigma_{k}}}+\underbrace{\nabla_{\mathbf{z}_{k}}\log{\left(\phi_{w}\left(\frac{\mathbf{z}_{k}-\sigma_{k}\epsilon^{*}_{\theta}(\mathbf{z}_{k},k)}{\alpha_{k}}\right)\right)}}_{\text{self guidance}},$ (20)
where recall that $\phi_{w}(\mathbf{z})=w$ simply extracts the weight component from $\mathbf{z}=[\mathbf{a},w]$.
Our method is referred to as Self-Weighted Guidance (SWG), since the DM samples from the target distribution $\mathbf{z}_{0}\sim p(\mathbf{z}_{0})$ by guiding itself, that is, the guidance signal is a function of the same model (see Algorithm 2). SWG provides an elegant way of performing exact guidance, since it computes gradients directly on the DM $\epsilon_{\theta}$ instead of an external guidance network with a dedicated —and possibly intricate— loss function (see Appendix C). This translates to not having to train a separate guidance network, thereby simplifying the training pipeline. Table 1 presents a conceptual comparison between SWG and other DM-based guidance approaches to offline RL.
Table 1: SWG in context: differences with other DM-based guidance approaches to offline RL.
Exact guidance Requires External Guidance Network
Diffuser janner2022diffuser No Yes
CEP cep Yes Yes
D-DICE diff_dice Yes Yes
SWG (ours) Yes No
@@ 4 Implementing SWG in offline RL
There are two main practical aspects regarding the implementation of SWG in the offline RL setup. First, recall that for notational and conceptual simplicity, SWG was presented in a way independent of the states. In practice, however, state-dependent DM’s are trained by considering the state as an input to the noise predictor network, that is, $\epsilon(\mathbf{z}_{k},k)=\epsilon(\mathbf{z}_{k},k,s)$ in Eq. (17). Second, SWG was introduced in Sec. 3 assuming a known weight function $w$, however, in offline RL the weights need to be estimated beforehand. We next present how to adapt and implement different weight formulations within the proposed SWG.
Expectile. The IDQL idql $\tau$-expectile weight is defined as $w_{\tau}^{2}(s,\mathbf{a})=\left|\tau-1_{\{Q(s,\mathbf{a})<V_{\tau}^{2}(s)\}}\right|$, where $V_{\tau}^{2}(s)$ is the value function learned with the expectile loss, and $1_{\{\cdot\}}$ is the indicator function. Since this weight takes only two possible values, $w_{\tau}^{2}(s,\mathbf{a})\in\{\tau,1-\tau\}\subset[0,1]$, it may fail to distinguish between similar actions by assigning them identical weights. To address this limitation, we propose a smooth variant of these weights, given by
@@
Where we apply a sigmoid function to the advantage $A(s,\mathbf{a})$ to obtain a continuous weight function. This weight function is utilized in our D4RL experiments (Sec.5.2), and the full training procedure for the value functions is detailed in Algorithm 3 in Appendix D.2.
Exponential. From the same value functions learned with expectile loss, we can define an exponential weight model derived from the IQL iql policy extraction method as $w(s,\mathbf{a})=e^{\beta A(s,\mathbf{a})}$, where $A(s,\mathbf{a})=Q(s,\mathbf{a})-V(s)$ is the advantage function. Though this formulation addresses the sparsity issue arising in the expectile weight, its exponential form could lead to arbitrarily large values, and is thus potentially unstable during training of SWG. Using the advantage function can help mitigate instabilities, as the value function serves as a normalizer for the Q-values.
Linex. Another exponential formulation can be obtained when learning the value function using the linex function $lx:u\mapsto\exp(u)-u$. As found by idql , this defines the following weight:
@@
where $V_{\text{exp}}(s)$ is the value function learned with the linex loss, and $\alpha$ is a temperature coefficient.
We emphasize that, among the various weighted behavior policies in the literature, we focus on avoiding querying out-of-sample actions, as they can lead to overestimation errors iql . Sec. 5.3 provides an experimental comparison of different weight formulations within SWG on D4RL, with further details in Appendix D.2.
@@
Lastly, in our implementation the strength of the guidance score is controlled via a hyperparameter $\rho$, called guidance scale diffusion_gans , as follows $\hat{\epsilon}:={\epsilon}_{\theta}(\mathbf{z}_{k},k)-\rho\sigma_{k}\nabla_{\mathbf{z}_{k}}\log{(\phi_{w}(\frac{\mathbf{z}_{k}-\sigma_{k}{\epsilon}_{\theta}(\mathbf{z}_{k},k)}{\alpha_{k}}))}$.
@@ 5 Experiments
The proposed SWG was implemented and validated through a set of experiments targeting the following questions (relevant experiments for each question in parentheses):
@@
    Can SWG sample from the target distribution? (5.1)
@@
    How does SWG perform against the state-of-the-art in the D4RL benchmark? (5.2)
@@
    How do different weight formulations affect SWG? (5.3)
@@
    How does SWG’s inference time scale with increasing model size? (5.4)
The experimental setup and benchmarks are in Appendix D, additional hyperparameter sensitivity experiments are in Appendices D.3 and E, and our code can be found in SWG repository.
@@ 5.1 Sampling from the target distribution
We considered the toy experiments from cep , which use a weight function $w(\mathbf{a})=\exp(-\beta\xi(\mathbf{a}))$, where $\xi(\mathbf{a})$ is an energy function, and $\beta$ is the inverse temperature coefficient. Figure 1 shows SWG samples from a spiral distribution and offers a qualitative assessment. The samples closely matching the ground truth for different values of $\beta$ suggest that SWG was able to sample from the target distribution, thus learning the relationship between the data distribution and the weights. The reader is referred to Appendix E for more toy examples of this kind.
Figure 1: SWG applied to a spiral distribution for different temperature coefficients $\beta$. Top: samples generated by SWG. Bottom: ground truth samples. When $\beta=0$, we have $w(\mathbf{a})=1$ and thus the target distribution coincides with the data distribution.
@@ 5.2 D4RL experiments
Using the D4RL benchmark fu2020d4rl , we compared SWG to classical offline RL methods (CQL cql and IQL iql ), diffusion-based resampling methods (SfBC sfbc and IDQL idql ), diffusion-guided methods (Diffuser janner2022diffuser , D-DICE (IGL) diff_dice , and GQPO (CEP)cep ), and to D-QL dql , which trains a DM using Q-values as regularization term.
Our experiments are performed across four task groups: 1) Locomotion is the main offline RL benchmark and primarily involves moving an agent across a plane. 2) Ant Maze is considered a more challenging task, where the goal is to guide an ant agent through a maze to reach a target. 3) Franka Kitchen involves goal-conditioned manipulation in a realistic kitchen environment. Lastly, 4) Adroit Pen focuses on high-dimensional dexterous manipulation to reorient a pen.
@@
These experiments were conducted with the expectile weight proposed in IDQL idql modified as detailed in Sec. 4, and using $K=15$ diffusion steps. See Appendix D.2 for more details.
@@
The results shown in Table 2 reveal that SWG performance was on par with state-of-the-art DM’s in challenging environments such as Ant Maze, Franka Kitchen and Adroit Pen tasks. In locomotion tasks, however, SWG achieved average performance with respect to previous methods. Nevertheless, the locomotion benchmark is considered to be saturated rafailovd5rl and does not necessarily represent a trustworthy criterion. Also note that SWG outperformed classical offline RL methods such as CQL and IQL in most tasks.
@@
Based on the D4RL results, we argue that SWG offers a tradeoff between having a simpler training setup with fewer components than current DM guidance-based approaches at the cost of a lower performance with respect to the top-performing method, D-DICE diff_dice .
@@
Table 2: SWG against the state-of-art on the D4RL benchmark: Mean and standard deviation of the normalized returns fu2020d4rl under $5$ random seeds. Values within $5\%$ of the best performing method in each task (row) are in bold font. See Appendix D for full experimental details.
Dataset Environment CQL IQL SfBC IDQL Diffuser D-QL D-DICE QGPO SWG (ours)
Medium-Expert HalfCheetah $62.4$ $86.7$ $\bf{92.6}$ $\bf{95.9}$ $79.8$ $\bf{96.8}$ $\bf{97.3}$ $\bf{93.5}$ $\bf{92.8}\pm 0.6$
Medium-Expert Hopper $98.7$ $91.5$ $\bf{108.6}$ $\bf{108.6}$ $\bf{107.2}$ $\bf{111.1}$ $\bf{112.2}$ $\bf{108.0}$ $\bf{109.9}\pm 3.2$
Medium-Expert Walker2d $\bf{111.0}$ $\bf{109.6}$ $\bf{109.8}$ $\bf{112.7}$ $\bf{108.4}$ $\bf{110.1}$ $\bf{114.1}$ $\bf{110.7}$ $\bf{110.7}\pm 1.7$
Medium HalfCheetah $44.4$ $47.4$ $45.9$ $51.0$ $44.2$ $51.1$ $\bf{60.0}$ $54.1$ $46.9\pm 0.2$
Medium Hopper $58.0$ $66.3$ $57.1$ $65.4$ $58.5$ $90.5$ $\bf{100.2}$ $\bf{98.0}$ $68.6\pm 4.4$
Medium Walker2d $79.2$ $78.3$ $77.9$ $82.5$ $79.7$ $\bf{87.0}$ $\bf{89.3}$ $\bf{86.0}$ $74.0\pm 2.5$
Medium-Replay HalfCheetah $46.2$ $44.2$ $37.1$ $45.9$ $42.2$ $\bf{47.8}$ $\bf{49.2}$ $\bf{47.6}$ $42.5\pm 1.5$
Medium-Replay Hopper $48.6$ $94.7$ $86.2$ $92.1$ $96.8$ $\bf{101.3}$ $\bf{102.3}$ $96.9$ $73.9\pm 2.3$
Medium-Replay Walker2d $26.7$ $73.9$ $65.1$ $85.1$ $61.2$ $\bf{95.5}$ $\bf{90.8}$ $84.4$ $60.6\pm 4.1$
Average (Locomotion) $63.9$ $76.9$ $75.6$ $82.1$ $75.3$ $\bf{88.0}$ $\bf{90.6}$ $\bf{86.6}$ $75.5$
Default AntMaze-umaze $74.0$ $87.5$ $92.0$ 94.0 - $\bf{93.4}$ $\bf{98.1}$ $\bf{96.4}$ $92.4\pm 2.6$
Diverse AntMaze-umaze $\bf{84.0}$ $62.2$ $\bf{85.3}$ 80.2 - $66.2$ $\bf{82.0}$ $74.4$ ${70}\pm 4.2$
Play AntMaze-medium $61.2$ $71.2$ $81.3$ 84.5 - $76.6$ $\bf{91.3}$ $83.6$ $\bf{89.6}\pm 2.3$
Diverse AntMaze-medium $53.7$ $70.0$ $82.0$ 84.8 - $78.6$ $\bf{85.7}$ $\bf{83.8}$ $\bf{88.2}\pm 5.2$
Play AntMaze-large $15.8$ $39.6$ $59.3$ 63.5 - $46.4$ $\bf{68.6}$ $\bf{66.6}$ $55.4\pm 4.6$
Diverse AntMaze-large $14.9$ $47.5$ $45.5$ 67.9 - $56.6$ $\bf{72.0}$ $64.8$ $55.2\pm 3$
Average (Ant Maze) $50.6$ $63.0$ $74.2$ 79.1 - $69.6$ $\bf{83}$ $78.3$ $75.1$
Human Adroit Pen $37.5$ $71.5$ $76$ $72.8$ $\bf{84.4}$ $72.8$ $78.5\pm 6.3$
Cloned Adroit Pen $39.2$ $37.3$ $64$ $57.3$ $\bf{83.8}$ $54.2$ $62.3\pm 6.7$
Average (Adroit Pen) $38.4$ $54.4$ $70$ $65.1$ $\bf{84.1}$ $63.5$ $70.4$
Partial Franka Kitchen $49.8$ $46.3$ $47.9$ $60.5$ $\bf{78.3}$ $69.0\pm 4.9$
Mixed Franka Kitchen $51.0$ $51.0$ $45.4$ $62.6$ $\bf{67.8}$ $61.0\pm 2.2$
Complete Franka Kitchen $43.8$ $62.5$ $77.9$ $\bf{84.0}$ $72.3\pm 2.6$
Average (Franka Kitchen) $48.2$ $53.3$ $57.1$ $\bf{69.0}$ $\bf{67.4}$

### v2

@@ 3.3 Augmented exact guidance
The required score to sample from the target distribution in equation 8 is given by
@@
The proof for Proposition 3.3 follows as an extension of (cep) to our setting (see Appendix A.1).
@@
Since the extended formulation in equation 8 is defined over $\mathbf{z}$, the gradients required to implement guidance in equation 10 are taken with respect to $\mathbf{z}_{k}$, rather than $\mathbf{a}_{k}$ (as in equation 7). This is a desired feature, since the actions and weights are highly coupled, and thus guiding them separately can lead to inefficient sampling.
The key insight of our proposal is that the intractable term in equation 10 can be recovered directly from the DM described above using the so-called data prediction formula (efron2011tweedie; variational_diffusion_models) given by:
@@
Using equation 11, and since the extraction function $\phi_{w}(\cdot)$ is linear, the target score $\nabla_{\mathbf{z}_{k}}\log p_{k}(\mathbf{z}_{k})$ in equation 10 can be computed with the proposed DM, as stated in Theorem 3.5 (proof in Appendix A.2).
@@
Let $\epsilon^{*}_{\theta}(\mathbf{z}_{k},k)$ be the DM found as the global minimizer of equation 9. The exact target score of equation 10 is given by:
$\nabla_{\mathbf{z}_{k}}\log p_{k}(\mathbf{z}_{k})=\underbrace{\nabla_{\mathbf{z}_{k}}\log q_{k}(\mathbf{z}_{k})}_{\frac{-\epsilon^{*}_{\theta}(\mathbf{z}_{k},k)}{\sigma_{k}}}+\underbrace{\nabla_{\mathbf{z}_{k}}\log{\left(\phi_{w}\left(\frac{\mathbf{z}_{k}-\sigma_{k}\epsilon^{*}_{\theta}(\mathbf{z}_{k},k)}{\alpha_{k}}\right)\right)}}_{\text{self guidance}},$
where recall that $\phi_{w}(\mathbf{z})=w$ simply extracts the weight component from $\mathbf{z}=[\mathbf{a},w]$ (Def. 3.1).
Our method is referred to as Self-Weighted Guidance (SWG), since the DM samples from the target distribution $\mathbf{z}_{0}\sim p(\mathbf{z}_{0})$ by guiding itself, that is, the guidance signal is a function of the same model (see Algorithm 2). SWG provides an elegant way of performing exact guidance, since it computes gradients directly on the DM $\epsilon_{\theta}$ instead of an external guidance network with a dedicated —and possibly intricate— loss function (see Sec. 4). This translates to not having to train a separate guidance network, thereby simplifying the training pipeline. Sec. 4 provides a theoretical comparison between SWG and other DM-based guidance approaches to offline RL, the results of which are summarized in Table 1.
@@
4 Relationship to previous guidance methods in offline RL
In this section, we provide an analysis of previous guidance methods alongside our own, highlighting theoretical differences and potential advantages. The two primary criteria for comparing previous methods with SWG are: i) whether the method provides exact guidance, and ii) whether it relies on an external guidance network, as summarized in Table 1.
Table 1: Summary of DM-based guidance approaches to offline RL.
Method Exact guidance Requires External Guidance Network Loss function
Diffuser (janner2022diffuser) No Yes Guidance network MSE over Q-values
CEP (cep) Yes Yes Contrastive loss (see equation 12)
D-DICE (diff_dice) Yes Yes In-sample loss (see equation 13)
SWG (ours) Yes No Diffusion loss (see equation 14)
Figure 1: Comparison between the proposed SWG approach (left) and the standard Diffusion Model with a separate Guidance Network (right). SWG jointly learns actions and weights within a single diffusion model, whereas standard approaches in diffusion for RL trains an additional guidance network to steer the diffusion model. The figure represents one forward pass to the diffusion model during training. Recall that $\mathbf{z_{0}}=[\mathbf{a_{0}},w_{0}]$.
Let us consider the two existing exact-guidance DM methods in the offline RL literature: CEP (cep) and D-DICE (diff_dice). Assuming a known weight function $w(\cdot)$, CEP learns the guidance network $g_{\theta}$ using the loss
where $k\sim\mathcal{U}(1,K)$, and for simplicity of notation we denote $g^{(i)}_{\theta,k}=g_{\theta}(s,\mathbf{a}_{k}^{(i)},k)$ and $\mathbf{a}_{0,k}^{(1:M)}=\{\mathbf{a}_{j}^{(i)},\text{s.t.},j\in\{0,k\},i\in\{1,\ldots,M\}\}$. Recall that $\mathbf{a}_{0}$ is the data and $\mathbf{a}_{k}$ are the noisy copies. Since equation 12 requires $M>1$ action samples from each state in the dataset, its evaluation is likely to require out-of-sample actions, thus possibly leading to overestimation errors (iql; diff_dice). CEP (cep) addresses this by pretraining a DM on the behavior policy and generating a synthetic action dataset containing $M>1$ samples for each state in the dataset.
In contrast, D-DICE proposes a loss that circumvents the need to query out-of-sample actions through
$\mathcal{L}_{D-DICE}(\theta)=\mathbb{E}_{k,\mathbf{a}_{0,k}}\Bigg[w(s,\mathbf{a}_{0})e^{-g_{\theta}(s,\mathbf{a}_{k},k)}+g_{\theta}(s,\mathbf{a}_{k},k)\Bigg],$ (13)
Although the objectives in equation 12 and equation 13 yield unbiased estimates of the intractable term in equation 5, their formulations introduce potential training instabilities arising from the exponential factor $e^{g_{\theta}}$. To address this, the authors of these loss functions proposed stabilization techniques: CEP employs self-normalization of the exponential term, whereas D-DICE alleviates the issue by clipping the exponential term. Furthermore, we contend that these loss functions are either lacking in direct interpretability, as in the case of D-DICE, or overly intricate, as in the case of CEP.
[…]
which is equivalent to maximizing a variational lower bound on the likelihood (variational_diffusion_models). That is, a simple MSE over predicted noise terms free from exponential terms prone to instabilities, thereby easing the method’s implementation. Another relevant feature of SWG is that it inherently defines an in-sample loss, thus addressing the out-of-sample problem with the CEP objective.
Also, since SWG produces action and weight samples $\mathbf{z}_{0}=[\mathbf{a}_{0},w_{0}]$ simultaneously, it is possible to define a resampling mechanism as described in Sec 2.2 to further improve performance. In Section 6, where our experimental results are presented, we include a variant of SWG called SWG-R, which uses resampling. These results demonstrate that SWG-R exhibits superior performance in comparison to the vanilla (non-resampling) SWG method.
@@ 5 Implementing SWG in offline RL
There are two main practical aspects regarding the implementation of SWG in the offline RL setup. First, recall that, for notational and conceptual simplicity, the SWG was presented in a way independent of the states. In practice, however, state-dependent DM’s are trained by considering the state as an input to the noise predictor network, that is, $\epsilon(\mathbf{z}_{k},k)=\epsilon(\mathbf{z}_{k},k,s)$ in equation 9. Second, SWG was introduced in Sec. 3 assuming a known weight function $w$, however, in offline RL the weights need to be estimated beforehand. We next present how to adapt and implement different weight formulations within the proposed SWG.
Expectile. The IDQL (idql) $\tau$-expectile weight is defined as $w_{\tau}^{2}(s,\mathbf{a})=\left|\tau-1_{\{Q(s,\mathbf{a})<V_{\tau}^{2}(s)\}}\right|$, where $V_{\tau}^{2}(s)$ is the value function learned with the expectile loss, and $1_{\{\cdot\}}$ is the indicator function. Since this weight takes only two possible values, $w_{\tau}^{2}(s,\mathbf{a})\in\{\tau,1-\tau\}\subset[0,1]$, it may fail to distinguish between similar actions by assigning them identical weights. To address this limitation, we propose a smooth variant of these weights, given by
@@
Here, we apply a sigmoid function to the advantage $A(s,\mathbf{a})$ to obtain a continuous weight function. This weight function is used in our D4RL experiments (Sec. 6.2), and the full training procedure for the value functions is detailed in Algorithm 3 in Appendix B.2.
Exponential. From the same value functions learned with expectile loss, we can define an exponential weight model derived from the IQL (iql) policy extraction method as $w(s,\mathbf{a})=e^{\beta A(s,\mathbf{a})}$, where $A(s,\mathbf{a})=Q(s,\mathbf{a})-V(s)$ is the advantage function. Though this formulation addresses the sparsity issue arising in the expectile weight, its exponential form could lead to arbitrarily large values, and is thus potentially unstable during training of SWG. Using the advantage function can help mitigate instabilities, as the value function serves as a normalizer for the Q-values.
Linex. Another exponential formulation can be obtained when learning the value function using the linex function $lx:u\mapsto\exp(u)-u$. As found by (idql), this defines the following weight:
@@
where $V_{\text{exp}}(s)$ is the value function learned by using the linex loss, and $\alpha$ is a temperature coefficient.
We emphasize that, among the various weighted behavior policies in the literature, we focus on avoiding querying out-of-sample actions, as they can lead to overestimation errors (iql). Sec. 6.3 provides an experimental comparison of different weight formulations within SWG on D4RL. More details on the implementation and training on the learning of $Q(s,a)$ and $V(s)$ are provided in Appendix B.2.
@@
In our implementation, we employ two techniques to refine the action generation process. First, we control the strength of the guidance score using a hyperparameter $\rho$, termed guidance scale (diffusion_gans), as follows $\hat{\epsilon}:={\epsilon}_{\theta}(\mathbf{z}_{k},k)-\rho\sigma_{k}\nabla_{\mathbf{z}_{k}}\log{(\phi_{w}(\frac{\mathbf{z}_{k}-\sigma_{k}{\epsilon}_{\theta}(\mathbf{z}_{k},k)}{\alpha_{k}}))}$. Second, we introduce SWG-R, a resampling variant of SWG, which resamples a candidate batch of $M$ actions, denoted as $\{a_{m}\}_{m=0}^{M}$, and selects the action $a^{*}$ associated with the maximum weight $w$.
@@ 6 Experiments
The proposed SWG was implemented and validated through a set of experiments targeting the following questions:
@@
    Can SWG sample from the target distribution? (Sec. 6.1)
@@
    How does SWG (with and without resampling) perform against the state-of-the-art in the D4RL benchmark? (Sec. 6.2)
@@
    How do different weight formulations affect SWG’s performance? (Sec. 6.3)
@@
    How does the guidance scale affect SWG’s performance? (Sec. 6.4)
  5. 5.
    How does SWG’s inference time scale with increasing model size? (Sec. 6.5)
Additional experimental details and benchmarks are given in Appendix B, with the full computational resources used in Appendix B.4. Our code is available at SWG repository.
@@ 6.1 Sampling from the target distribution
We considered the toy experiments from (cep), which use a weight function $w(\mathbf{a})=\exp(-\beta\xi(\mathbf{a}))$, where $\xi(\mathbf{a})$ is an energy function, and $\beta$ is the inverse temperature coefficient. Figures 2 and 3 show SWG samples from a spiral and an 8-Gaussian-mixture distributions, respectively, and offer a qualitative assessment of its performance. In both cases, samples closely matching the ground truth for different values of $\beta$ suggest that SWG was able to sample from the target distributions, thus learning the relationship between the data distribution and the weights in both cases. The reader is referred to the Appendix B.1 for further details and more examples with different target distributions.
Figure 2: SWG applied to a spiral distribution for different temperature coefficients $\beta$. Top: samples generated by SWG. Bottom: ground truth samples. When $\beta=0$, we have $w(\mathbf{a})=1$ and thus the target distribution matches the data distribution.
Figure 3: SWG applied to an 8-Gaussian mixture distribution for different temperature coefficients $\beta$. Top: samples generated by SWG. Bottom: ground truth samples. When $\beta=0$, we have $w(\mathbf{a})=1$ and thus the target distribution matches the data distribution.
@@ 6.2 D4RL experiments
Using the D4RL benchmark (fu2020d4rl), we implemented two variants of the proposed method: SWG as described above, and SWG-R, which includes action resampling on top of our method. We considered the following methods as benchmarks:
  • Classical offline RL: CQL (cql) and IQL (iql)
  • Diffusion-based resampling: SfBC (sfbc) and IDQL (idql)
  • Diffusion-guided: Diffuser (janner2022diffuser), D-DICE (diff_dice), and QGPO (CEP) (cep)
  • D-QL (dql), which trains a DM using Q-values as regularization term
Furthermore, our experiments are performed across four task groups:
  1. 1.
    Locomotion: the main offline RL benchmark, which features moving an agent across a plane.
  2. 2.
    Ant Maze: a more challenging task, where an ant agent is guided through a maze to reach a target.
  3. 3.
    Franka Kitchen: involves goal-conditioned manipulation in a realistic kitchen environment.
  4. 4.
    Adroit Pen: focuses on high-dimensional dexterous manipulation to reorient a pen.
@@
These experiments were conducted with the expectile weight proposed in IDQL (idql) modified as detailed in Sec. 5, and using $K=15$ diffusion steps. See Appendix B.2 for more details.
@@
The results shown in Table 2 reveal that the resampled version of our proposal SWG-R provided competitive performance over state-of-the-art DMs in the challenging environment tasks from Franka Kitchen and Adroit Pen. In Ant Maze, SWG-R also reported very competitive figures, outperforming all methods except D-DICE. However, in locomotion tasks, both SWG and SWG-R achieved average performance compared to previous methods. Also, note that SWG outperformed classical offline RL methods such as CQL and IQL in most tasks.
@@
Based on the D4RL results, we argue that SWG-R offers a tradeoff between having a simpler training setup with fewer components than current DM guidance-based approaches at the cost of a lower performance with respect to the top-performing method, D-DICE (diff_dice), for the particular case of the Locomotion dataset. Moreover, incorporating resampling was pivotal for the enhanced performance of SWG-R over SWG. This addition strengthened the effectiveness of our framework, similarly to previous approaches. However, it is worth noting that the performance gain of SWG-R comes at the cost of an expanded hyperparameter search space during inference (see Appendix B.2). Additionally, we remark that the resampling was performed using the separate weight model $w$ (i.e., expectile weights).
@@
Table 2: SWG against the state-of-art on the D4RL benchmark: Mean and standard deviation of the normalized returns (fu2020d4rl) under 55 random seeds. Values within $5\%$ of the best performing method in each task (row) are in bold font. SWG and SWG-R correspond to the standard and resampling version of our method, respectively. We used acronyms ME: Medium-Expert, M: Medium, MR: Medium-Replay, Avg: Average, HC: HalfCheetah, Hopp: Hopper, W2D: Walker2D, Kitchen: Franka Kitchen, Pen: Adroit Pen. See Appendix B.2 for full experimental details.
Dataset Env CQL IQL SfBC IDQL Diffuser D-QL D-DICE QGPO SWG SWG-R
ME HC $62.4$ $86.7$ $\bf{92.6}$ $\bf{95.9}$ $79.8$ $\bf{96.8}$ $\bf{97.3}$ $\bf{93.5}$ $\bf{92.8}\pm 0.6$ $\bf{93.5}\pm 0.5$
ME Hopp $98.7$ $91.5$ $\bf{108.6}$ $\bf{108.6}$ $\bf{107.2}$ $\bf{111.1}$ $\bf{112.2}$ $\bf{108.0}$ $\bf{109.9}\pm 3.2$ $\bf{111.0}\pm 0.2$
ME W2D $\bf{111.0}$ $\bf{109.6}$ $\bf{109.8}$ $\bf{112.7}$ $\bf{108.4}$ $\bf{110.1}$ $\bf{114.1}$ $\bf{110.7}$ $\bf{110.7}\pm 1.7$ $\bf{112.5}\pm 0.3$
M HC $44.4$ $47.4$ $45.9$ $51.0$ $44.2$ $51.1$ $\bf{60.0}$ $54.1$ $46.9\pm 0.2$ $46.9\pm 0.2$
M Hopp $58.0$ $66.3$ $57.1$ $65.4$ $58.5$ $90.5$ $\bf{100.2}$ $\bf{98.0}$ $68.6\pm 4.4$ $74.6\pm 5.2$
M W2D $79.2$ $78.3$ $77.9$ $82.5$ $79.7$ $\bf{87.0}$ $\bf{89.3}$ $\bf{86.0}$ $74.0\pm 2.5$ $74.8\pm 2.1$
MR HC $46.2$ $44.2$ $37.1$ $45.9$ $42.2$ $\bf{47.8}$ $\bf{49.2}$ $\bf{47.6}$ $42.5\pm 1.5$ $42.5\pm 1.5$
MR Hopp $48.6$ $94.7$ $86.2$ $92.1$ $96.8$ $\bf{101.3}$ $\bf{102.3}$ $96.9$ $73.9\pm 2.3$ $73.9\pm 2.3$
MR W2D $26.7$ $73.9$ $65.1$ $85.1$ $61.2$ $\bf{95.5}$ $\bf{90.8}$ $84.4$ $60.6\pm 4.1$ $79.9\pm 4.5$
Avg (Loc) $63.9$ $76.9$ $75.6$ $82.1$ $75.3$ $\bf{88.0}$ $\bf{90.6}$ $\bf{86.6}$ $75.5$ 78.8
Default Umaze $74.0$ $87.5$ $92.0$ $\bf{94.0}$ - $\bf{93.4}$ $\bf{98.1}$ $\bf{96.4}$ $92.4\pm 2.6$ $\bf{98.3}\pm 1.7$
Diverse Umaze $\bf{84.0}$ $62.2$ $\bf{85.3}$ 80.2 - $66.2$ $\bf{82.0}$ $74.4$ ${70}\pm 4.2$ $71.7\pm 8.3$
Play Med $61.2$ $71.2$ $81.3$ 84.5 - $76.6$ $\bf{91.3}$ $83.6$ $\bf{89.6}\pm 2.3$ $\bf{90.0}\pm 2.9$
Diverse Med $53.7$ $70.0$ $82.0$ 84.8 - $78.6$ 85.7{85.7} 83.8{83.8} $\bf{88.2}\pm 5.2$ $\bf{91.7}\pm 1.7$
Play Large $15.8$ $39.6$ $59.3$ 63.5 - $46.4$ $\bf{68.6}$ $\bf{66.6}$ $55.4\pm 4.6$ $\bf{66.7}\pm 3.3$
Diverse Large $14.9$ $47.5$ $45.5$ 67.9 - $56.6$ $\bf{72.0}$ $64.8$ $55.2\pm 3$ $63.3\pm 8.3$
Avg (AntMaze) $50.6$ $63.0$ $74.2$ 79.1 - $69.6$ $\bf{83}$ $78.3$ $75.1$ 80.3
Human Pen $37.5$ $71.5$ - $76$ - $72.8$ 84.4{84.4} $72.8$ $78.5\pm 6.3$ $\bf{103.0}\pm 5.0$
Cloned Pen $39.2$ $37.3$ - $64$ - $57.3$ 83.8{83.8} $54.2$ $62.3\pm 6.7$ $\bf{91.7}\pm 10.5$
Avg (Adroit Pen) $38.4$ $54.4$ - $70$ - $65.1$ 84.1{84.1} $63.5$ $70.4$ 97.4
Partial Kitchen $49.8$ $46.3$ $47.9$ - - $60.5$ $\bf{78.3}$ - $69.0\pm 4.9$ $73.3\pm 1.4$
Mixed Kitchen $51.0$ $51.0$ $45.4$ - - $62.6$ $\bf{67.8}$ - $61.0\pm 2.2$ $\bf{65.0}\pm 0.8$
Complete Kitchen $43.8$ $62.5$ $77.9$ - - $\bf{84.0}$ - - $72.3\pm 2.6$ $72.3\pm 2.6$
Avg (Kitchen) $48.2$ $53.3$ $57.1$ - - $\bf{69.0}$ - - $\bf{67.4}$ 70.2